OpenStudy (anonymous):

(x)^2+(x+7)^2=(13)^2 this is the set up for this word problem...... The length of one leg of a right triangle is seven inches more than the length of the other leg @Mchilds15

OpenStudy (mchilds15):

Do I need to solve it for you?

OpenStudy (anonymous):

I don't know where to go from there i set it up but i don't know how to solve it

OpenStudy (mchilds15):

I see. Okay so, we're solving for x, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its using the Pythagorean theorem

OpenStudy (mchilds15):

Sooooooo basically you have x^2 + x^2 + 49= 169

OpenStudy (anonymous):

what do i do with the (x+7)^2

OpenStudy (anonymous):

basically the problem is xsquared + x+7squared=13squared

OpenStudy (mchilds15):

Lol the x^2+49 is the (x+7)^2

OpenStudy (anonymous):

ok

OpenStudy (mchilds15):

My equation is the simplified version. Anywho, now that you have x^2 + x^2 + 49= 169, you need to simplify it So, what are you going to do first?

OpenStudy (anonymous):

do I combine the like terms which is the xsquared and xsquared??

OpenStudy (mchilds15):

Hm..... I'm second guessing myself... But yeah, do that, then subtract the 49 from 169. (I haven't done this in years, so sorry if this ends up wrong. I'm trying to remember)

OpenStudy (anonymous):

so when u ccombine those wont it be 2xsquared??

OpenStudy (mchilds15):

Yep

OpenStudy (anonymous):

thank u so much im not really good at math so I been trying to find online help and u saved my life today thank u again and if I need help again can I just tag u in a question??

OpenStudy (mchilds15):

Yeah lol So did you solve for x?

OpenStudy (anonymous):

yea I got it thank u